Market price is $5620Y2 2011 Retained earnings, January 1 $ 1,074,900 $ 1,267,700 292,000 Net income 220,200 Total $ 1,559,700 $ 1,295,100 DividMul Sllall INC. Comparative Income Statement For the Years Ended December 31, 20Y2 and 20Y1 20Y2 20Y1 es $ 1,768,425 $ 1,629,Current assets Cash $ 238,980 Marketable securities 361,710 321,200 $ 322,390 534,240 299,300 Accounts receivable (net) Inven11. Asset turnover 12. Return on total assets 0.1 x % 13. Return on stockholders equity 0.2 x % 14. Return on common stockho
Answer #1
11 Asset turnover:-
a. Net Sales $ 1,768,425.00
b. Average Total assets
opening + closing 3439539+ 4051370 = $ 3,745,454.50
2 2
Ratio (a/b) 0.5 times
12 Return on Total Assets Ratio:-  
a. Net income $         292,000
b. Interest Expense (1-tax rate) (Tax rate = (39800/331800)*100 = 12%) 120000 (1-0.12) $         105,600
c Net income + Interest Expense (1-tax rate) $         397,600
d Average Total assets
opening + closing 3439539+ 4051370 = $ 3,745,454.50
2 2
Ratio (c/d *100) 10.6%
13. Return on stockholders' equity:-
a. Net income $         292,000
b. Average stockholders' equity
opening + closing 1867700 + 2132300 = $      2,000,000
2 2
Ratio (a/b *100) 14.6%
14. Return on common stockholders' equity:-
a. Earnings available for Equity shareholders = Net Income - Preferred Dividends = 292000-9800 = $         282,200
b. Average common stockholder equity
Common Stockholders' equity= Total Stockholders' equity -Preferred stock
Opening Common Stockholders' equity = 1867700-280000 = 1,587,700
Closing Common Stockholders' equity = 2132300-280000 = 1,852,300
opening + closing 1587700+ 1852300 = $      1,720,000
2 2
Ratio (a/b *100) 16.4%
15. Earnings per share on common stock:-
a. Earnings available for Equity shareholders = Net Income - Preferred Dividends = 292000-9800 = $         282,200
b. Average Number of shares of common stock
opening + closing 32,000 + 32,000 =               32,000
2 2
Ratio (a/b) $             8.8
16. Price earnings Ratio on Common Stock :-  
a. Market price per share $             56.00
b. Earnings per share $                 8.8
Ratio (a/b) 6.35 times
17. Dividends per share of Common Stock :-  
a. Common Stock Dividend declared $           17,600
b. Number of shares of common stock               32,000
Ratio (a/b) $           0.55
18 Dividend Yield on Common Stock Ratio:-
a. Dividend Per Share $               0.55
b. Market Price per share $             56.00
Dividend Yield Ratio (a/b*100) 1.0%

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.